The stemplot along with some summary statistics of the grades of a statistics class are given below. How many of these scores are identified as MILD outliers? Stem-and-Leaf Display: grade Stem-and-leaf of grade N 90 Leaf Unit 1.0 6 0001122233344 6 68899 7 0111233 7 667 8 69 9 Descriptive Statistics: grade Variable N Q1 03 grade 90 53 63 O 2
Answer #1

here IQR=Q3-Q1 =63-53 =10

lower mild fence =Q1-1.5*IQR=53-1.5*10=38

upper mild fence =Q3+1.5*IQR=63+1.5*10=78

lower extreme fence =Q1-3*IQR=53-3*10=23

upper  extreme fence =Q3+3*IQR=63+3*10=93

therefore mild outliers will fall between 23 and 38 as well between 78 and 93

number of values in mild outlier range are =86 and 89

therefore correct answer is 2